# Invariant subspaces and PCA in nearly matrix multiplication time

### v3

@@ E.3 Singular values, singular gaps, and condition number
$O\left(\log(n)\log(\tfrac{1}{\delta\sigma_{\min}(\boldsymbol{\mathrm{A}}^{\prime})})\log^{3}(\tfrac{n}{\delta\sigma_{\min}(\boldsymbol{\mathrm{A}}^{\prime})})\right)=O\left(\log(n)\log(\tfrac{\kappa(\boldsymbol{\mathrm{A}})}{\delta})\log^{3}(\tfrac{n\kappa(\boldsymbol{\mathrm{A}})}{\delta})\right)$
bits. Then we can set $\widetilde{\kappa}=\frac{1}{\widetilde{\sigma}_{\min}^{\prime}}\in\left[\kappa(\boldsymbol{\mathrm{A}}),32\kappa(\boldsymbol{\mathrm{A}})\right]$. The total success probability is $1-\tfrac{2\delta}{3}-\tfrac{\delta}{3}=1-\delta$. ∎
@@
$O\left(T_{\mathsf{MM}}(n)\left(\log(\tfrac{n}{\operatorname{gap}_{k}})\log(\tfrac{1}{\operatorname{gap}_{k}})+\log(n\kappa(\boldsymbol{\mathrm{S}}))\log(\kappa(\boldsymbol{\mathrm{S}})+\log\left(\log(\tfrac{\kappa(\boldsymbol{\mathrm{S}})}{\epsilon\operatorname{gap}_{k}})\right)\right)\right)$
@@
$O\left(\log(n)\log^{4}(\tfrac{n}{\operatorname{gap}_{k}})+\log(\kappa(\boldsymbol{\mathrm{S}})\right)$
@@
Algorithm 5 $\mathsf{PROJECTOR}$.
1:
$\mathsf{PROJECTOR}$
8:return $\widetilde{\boldsymbol{\mathrm{\Pi}}}$.
@@ Appendix F DFT background
Density Functional Theory (DFT) [KS65], which was awarded the Nobel prize in Chemistry in 1998, is considered as one of the most common methods to perform electronic structure calculations thanks to its ab initio character. That means that no input besides the initial atomic structure is required to determine the relaxed geometry, energy levels, and the electron structure of a given material, in their bulk or nanostructured form. The main idea behind DFT is to describe the system and its properties by the electron density only. DFT calculations are widely used in industry and in academia to predict, for example, the properties of novel materials [MFR+19] or to optimize the performance of batteries [HYLZ19], solar cells [XY17], or nanoelectronic devices [KSS+20]. Many scientific libraries implementing DFT algorithms have been developed, which persistently occupy supercomputing clusters and receive up to tens of thousands of citations annually at the time of this writing [KF96, GAB+17, SAG+02, HISV14]. Despite the success of DFT calculations, important theoretical aspects of their underlying algorithms remain unclear.
@@
where $E_{F}$ is the Fermi level of the system, $E_{i}$ is the energy level (eigenvalue) corresponding to the wave function $\psi_{i}(\boldsymbol{\mathrm{r}})$, and $f(E_{i};E_{F})$ is the occupation term. At zero temperature $(T=0K)$, $f(E_{i};E_{F})$ is equal to one for occupied states ($E_{i}<E_{F}$) and zero otherwise. [^3] The definition of $f(E;E_{F})$ can vary to allow for partial occupations, i.e. at nonzero temperature $T>0$ it can be replaced by the Fermi-Dirac distribution $1-\frac{1}{1+\exp((E_{i}-E_{F})/k_{B}T)}$, where $k_{B}$ is Boltzmann’s constant. This case is not considered in this work. The $E_{F}$, $E_{i}$, and $\psi_{i}$ are the unknown quantities for which there exists no analytical solution, except for some special cases. Approximate numerical solutions can be sought after expanding the wave functions into a suitable basis. In this case, the $\psi_{i}$’s are approximated by a set of $n=\Theta(n_{a})$ basis functions $\chi_{j}(\boldsymbol{\mathrm{r}}),j\in[n]$, typically localized around atomic positions. [^4] A common alternative is to approximate the electronic wave function using Plane Waves. In this work we focus on localized basis sets. Common basis sets that lead to efficient algorithms to solve the eigenvalue problem of Eq. (20) include Linear Combination of Atomic Orbitals (LCAO) [CK99], Gaussian-Type Orbitals (GTO) [BHM69], or Maximally Localized Wannier Functions (MLWF) [MMY+12]. In this context, the wave functions take the form of linear combinations of the $n$ localized orbitals. In particular, for all $i\in[n]$
@@
which is equal to the identity matrix $\boldsymbol{\mathrm{I}}$ if the basis elements are orthogonal (e.g. $\boldsymbol{\mathrm{S}}=\boldsymbol{\mathrm{I}}$ for MLWF). The entries of $\boldsymbol{\mathrm{H}}$ and $\boldsymbol{\mathrm{S}}$ represent the interaction between two basis elements that can be located on the same or on different atoms. Furthermore, due to the complex conjugate property of the inner product, both $\boldsymbol{\mathrm{H}}$ and $\boldsymbol{\mathrm{S}}$ are Hermitian matrices. Moreover, $\boldsymbol{\mathrm{S}}$ is positive-definite ([May03, Chapter 3]). Putting everything together, the expansion coefficients $c_{ij}$ are obtained by solving the following generalized eigenvalue problem as already formulated in Eq. (1):
@@ F.1 Linear scaling methods
Because of the strong localization of the $\chi_{i}(\boldsymbol{\mathrm{r}})$’s, the $(i,j)$ interactions rapidly decay with the distance between the atoms and can be discarded beyond a pre-defined cut-off radius. As a consequence, $\boldsymbol{\mathrm{H}}$ and $\boldsymbol{\mathrm{S}}$ are often banded matrices with sometimes a sparse band. It should also be noted that at equilibrium, the number of negatively charged electrons present in the system (which corresponds to the integral of $n(\boldsymbol{\mathrm{r}})$ over space) must exactly compensate the number of positively charged protons $Z_{p}$. This allows for the computation of the Fermi level: from Eq. (25) and the definition of $\boldsymbol{\mathrm{S}}$, it follows that $\operatorname{trace}(\boldsymbol{\mathrm{P}}\boldsymbol{\mathrm{S}})=Z_{p}$, which means that one can perform a binary search on the Fermi level $E_{F}$ until $\operatorname{trace}(\boldsymbol{\mathrm{P}}\boldsymbol{\mathrm{S}})=Z_{p}$ is satisfied. These observation has led to the development of so-called linear scaling methods, which aim to exploit the sparsity of the matrices to obtain faster solutions [Yan91, GP92, Koh93, Koh96, Gal96, Goe99, GS03]. Typically, the density matrix and the Fermi level are iteratively computed, until all required constraints are satisfied, i.e., the density matrix should be idempotent and $\operatorname{trace}(\boldsymbol{\mathrm{P}}\boldsymbol{\mathrm{S}})=Z_{p}$. This can be done, for example, by approximating the matrix sign function using a Newton iteration [VBH12]. To optimize performance, heuristics such as truncating matrix elements with negligible magnitude are applied, at the cost of decreasing the solution accuracy. Empirical evidence shows that such methods tend to scale nearly linearly to the system size, however, they do not exhibit provable theoretical guarantees. Similar methods have been studied based on finite-differences, which aim to approximate the Fermi distribution with polynomial expansions [ZSTC06b, ZSTC06a]. Finally, the work of [GE95, VXCB16] offer the possibility to achieve lower complexities than $O(n^{3})$ for both the eigenvalues and eigenvectors of structured matrices, by leveraging the fast multipole method. However, an end-to-end stability analysis of those algorithms remains open. Indeed, we are not aware of any end-to-end analysis with provable approximation guarantees that has lower than $O(n^{3})$ worst-case complexity for any of the aforementioned algorithms.
@@ F.3 Electron density queries
First we recall that $\epsilon\in(0,1/12)$ and note that $\|\widetilde{\boldsymbol{\mathrm{P}}}\|\leq\|\boldsymbol{\mathrm{P}}\|+\epsilon\leq 1+\epsilon.$ The quadratic form $\boldsymbol{\mathrm{x}}^{*}\widetilde{\boldsymbol{\mathrm{P}}}\boldsymbol{\mathrm{x}}$ is computed in two steps: $\boldsymbol{\mathrm{y}}\leftarrow\boldsymbol{\mathsf{fl}}(\widetilde{\boldsymbol{\mathrm{P}}}\boldsymbol{\mathrm{x}})$ and $\widetilde{n}(\boldsymbol{\mathrm{r}})=\boldsymbol{\mathsf{fl}}(\boldsymbol{\mathrm{x}}^{*}\boldsymbol{\mathrm{y}})$. From [Hig02], Eq. (3.12) and Lemma 6.6, we know that the matrix-vector product satisfies $\boldsymbol{\mathrm{y}}=\widetilde{\boldsymbol{\mathrm{P}}}\boldsymbol{\mathrm{x}}+\boldsymbol{\mathrm{e}}$ where $\|\boldsymbol{\mathrm{e}}\|=\|\widetilde{\boldsymbol{\mathrm{P}}}\boldsymbol{\mathrm{x}}-\boldsymbol{\mathrm{y}}\|\leq\sqrt{n}\eta_{n}\|\widetilde{\boldsymbol{\mathrm{P}}}\|\|\boldsymbol{\mathrm{x}}\|$, and $\eta_{n}:=\frac{n\textbf{{u}}}{1-n\textbf{{u}}}$. For the chosen u we have $\sqrt{n}\eta_{n}\leq 2\epsilon$. From the same equations, the subsequent dot product satisfies $\widetilde{n}(\boldsymbol{\mathrm{r}})=\boldsymbol{\mathsf{fl}}(\boldsymbol{\mathrm{x}}^{*}\boldsymbol{\mathrm{y}})=\boldsymbol{\mathrm{x}}^{*}\boldsymbol{\mathrm{y}}+\epsilon$ where $|\epsilon|\leq\eta_{n}\|\boldsymbol{\mathrm{x}}\|\|\boldsymbol{\mathrm{y}}\|\leq\eta_{n}\|\boldsymbol{\mathrm{x}}\|\left(\|\widetilde{\boldsymbol{\mathrm{P}}}\boldsymbol{\mathrm{x}}\|+\|\boldsymbol{\mathrm{e}}\|\right)\leq\eta_{n}\|\widetilde{\boldsymbol{\mathrm{P}}}\|\|\boldsymbol{\mathrm{x}}\|^{2}\left(1+\sqrt{n}\eta_{n}\right)\leq 4\epsilon\|\boldsymbol{\mathrm{x}}\|^{2}$. Then $|n(\boldsymbol{\mathrm{r}})-\widetilde{n}(\boldsymbol{\mathrm{r}})|=|\boldsymbol{\mathrm{x}}^{*}\boldsymbol{\mathrm{P}}\boldsymbol{\mathrm{x}}-(\boldsymbol{\mathrm{x}}^{*}\boldsymbol{\mathrm{y}}+\epsilon)|=|\boldsymbol{\mathrm{x}}^{*}\boldsymbol{\mathrm{P}}\boldsymbol{\mathrm{x}}-\boldsymbol{\mathrm{x}}^{*}\widetilde{\boldsymbol{\mathrm{P}}}\boldsymbol{\mathrm{x}}-\boldsymbol{\mathrm{x}}^{*}\boldsymbol{\mathrm{e}}-\epsilon|\leq|\boldsymbol{\mathrm{x}}^{*}\boldsymbol{\mathrm{P}}\boldsymbol{\mathrm{x}}-\boldsymbol{\mathrm{x}}^{*}\widetilde{\boldsymbol{\mathrm{P}}}\boldsymbol{\mathrm{x}}|+|\boldsymbol{\mathrm{x}}^{*}\boldsymbol{\mathrm{e}}|+|\epsilon|\leq 2\epsilon\|\boldsymbol{\mathrm{x}}\|^{2}+2\epsilon\|\boldsymbol{\mathrm{x}}\|^{2}+4\epsilon\|\boldsymbol{\mathrm{x}}\|^{2}=8\epsilon\|\boldsymbol{\mathrm{x}}\|^{2}.$
@@
If $N$ is asymptotically larger than the system size, i.e., $N=O(\operatorname{poly}(n))$, the query complexity can be reduced by applying the Johnson-Lindenstrauss (JL) lemma [JL84], at the cost of a polynomial dependence on the accuracy. In brief, since the matrix $\boldsymbol{\mathrm{P}}\boldsymbol{\mathrm{S}}$ is an orthogonal projector, and therefore idempotent, it can be deduced that
@@
where $\boldsymbol{\mathrm{L}}$ is the lower triangular Cholesky factor of $\boldsymbol{\mathrm{S}}^{-1}$. Then $\|\boldsymbol{\mathrm{x}}^{*}\boldsymbol{\mathrm{P}}\boldsymbol{\mathrm{L}}^{-*}\|^{2}$ can be approximated by $\|\boldsymbol{\mathrm{x}}^{*}\boldsymbol{\mathrm{P}}\boldsymbol{\mathrm{L}}^{-*}\boldsymbol{\mathrm{T}}\|^{2}$, where $\boldsymbol{\mathrm{T}}$ is a $n\times r$ sparse random sign matrix scaled by $1/\sqrt{r}$ [Ach01], with $r=O(\log(N)/\epsilon^{2})$ columns. The representation of $\boldsymbol{\mathrm{T}}$ requires $O(n\times r)=O(n\log(N)/\epsilon^{2})$ random bits. Alternative random matrix distributions satisfying the JL property, with varying sparsity, runtime, and dimension requirements exist in the literature [AC06, KN14, NN13]. The value of $r$ is worst-case optimal [LN17], however, tighter bounds have been studied in the literature [BRS11, SL22, BN22, ETW24], thus potentially leading to more practical algorithms. The full analysis is omitted.

### v4

@@ E.3 Singular values, singular gaps, and condition number
$O\left(\log(n)\log^{4}(\tfrac{n}{\delta\sigma_{\min}(\boldsymbol{\mathrm{A}}^{\prime})})\right)=O\left(\log(n)\log^{4}(\tfrac{n\kappa(\boldsymbol{\mathrm{A}})}{\delta})\right)$
bits. Then we can set $\widetilde{\kappa}=\frac{1}{\widetilde{\sigma}_{\min}^{\prime}}\in\left[\kappa(\boldsymbol{\mathrm{A}}),32\kappa(\boldsymbol{\mathrm{A}})\right]$. The success probability is $1-\tfrac{2\delta}{3}-\tfrac{\delta}{3}=1-\delta$. ∎
@@
$O\left(\log(n)\log^{4}(\tfrac{n}{\operatorname{gap}_{k}})+\log(\kappa(\boldsymbol{\mathrm{S}}))\right)$
@@ Appendix F DFT background
Density Functional Theory (DFT) [88], which was awarded the Nobel prize in Chemistry in 1998, is considered as one of the most common methods to perform electronic structure calculations thanks to its ab initio character. That means that no input besides the initial atomic structure is required to determine the relaxed geometry, energy levels, and the electron structure of a given material, in their bulk or nanostructured form. The main idea behind DFT is to describe the system and its properties by the electron density only. DFT calculations are widely used in industry and in academia to predict, for example, the properties of novel materials [102] or to optimize the performance of batteries [69], solar cells [142], or nanoelectronic devices [85]. Many scientific libraries implementing DFT algorithms have been developed, which persistently occupy supercomputing clusters and receive up to tens of thousands of citations annually at the time of this writing [89, 59, 127, 74]. Despite the success of DFT calculations, important theoretical aspects of their underlying algorithms remain unclear.
@@
where $E_{F}$ is the Fermi level of the system, $E_{i}$ is the energy level (eigenvalue) corresponding to the wave function $\psi_{i}(\boldsymbol{\mathrm{r}})$, and $f(E_{i};E_{F})$ is the occupation term. At zero temperature $(T=0K)$, $f(E_{i};E_{F})$ is equal to one for occupied states ($E_{i}<E_{F}$) and zero otherwise. [^3] The definition of $f(E;E_{F})$ can vary to allow for partial occupations, i.e. at nonzero temperature $T>0$ it can be replaced by the Fermi-Dirac distribution $1-\frac{1}{1+\exp((E_{i}-E_{F})/k_{B}T)}$, where $k_{B}$ is Boltzmann’s constant. This case is not considered in this work. The $E_{F}$, $E_{i}$, and $\psi_{i}$ are the unknown quantities for which there exists no analytical solution, except for some special cases. Approximate numerical solutions can be sought after expanding the wave functions into a suitable basis. In this case, the $\psi_{i}$’s are approximated by a set of $n=\Theta(n_{a})$ basis functions $\chi_{j}(\boldsymbol{\mathrm{r}}),j\in[n]$, typically localized around atomic positions. [^4] A common alternative is to approximate the electronic wave function using Plane Waves. In this work we focus on localized basis sets. Common basis sets that lead to efficient algorithms to solve the eigenvalue problem of Eq. (20) include Linear Combination of Atomic Orbitals (LCAO) [32], Gaussian-Type Orbitals (GTO) [22], or Maximally Localized Wannier Functions (MLWF) [101]. In this context, the wave functions take the form of linear combinations of the $n$ localized orbitals. In particular, for all $i\in[n]$
@@
which is equal to the identity matrix $\boldsymbol{\mathrm{I}}$ if the basis elements are orthogonal (e.g. $\boldsymbol{\mathrm{S}}=\boldsymbol{\mathrm{I}}$ for MLWF). The entries of $\boldsymbol{\mathrm{H}}$ and $\boldsymbol{\mathrm{S}}$ represent the interaction between two basis elements that can be located on the same or on different atoms. Furthermore, due to the complex conjugate property of the inner product, both $\boldsymbol{\mathrm{H}}$ and $\boldsymbol{\mathrm{S}}$ are Hermitian matrices. Moreover, $\boldsymbol{\mathrm{S}}$ is positive-definite ([103, Chapter 3]). Putting everything together, the expansion coefficients $c_{ij}$ are obtained by solving the following generalized eigenvalue problem as already formulated in Eq. (1):
@@ F.1 Linear scaling methods
Because of the strong localization of the $\chi_{i}(\boldsymbol{\mathrm{r}})$’s, the $(i,j)$ interactions rapidly decay with the distance between the atoms and can be discarded beyond a pre-defined cut-off radius. As a consequence, $\boldsymbol{\mathrm{H}}$ and $\boldsymbol{\mathrm{S}}$ are often banded matrices with sometimes a sparse band. It should also be noted that at equilibrium, the number of negatively charged electrons present in the system (which corresponds to the integral of $n(\boldsymbol{\mathrm{r}})$ over space) must exactly compensate the number of positively charged protons $Z_{p}$. This allows for the computation of the Fermi level: from Eq. (25) and the definition of $\boldsymbol{\mathrm{S}}$, it follows that $\operatorname{trace}(\boldsymbol{\mathrm{P}}\boldsymbol{\mathrm{S}})=Z_{p}$, which means that one can perform a binary search on the Fermi level $E_{F}$ until $\operatorname{trace}(\boldsymbol{\mathrm{P}}\boldsymbol{\mathrm{S}})=Z_{p}$ is satisfied. These observation has led to the development of so-called linear scaling methods, which aim to exploit the sparsity of the matrices to obtain faster solutions [143, 57, 86, 87, 56, 62, 63]. Typically, the density matrix and the Fermi level are iteratively computed, until all required constraints are satisfied, i.e., the density matrix should be idempotent and $\operatorname{trace}(\boldsymbol{\mathrm{P}}\boldsymbol{\mathrm{S}})=Z_{p}$. This can be done, for example, by approximating the matrix sign function using a Newton iteration [136]. To optimize performance, heuristics such as truncating matrix elements with negligible magnitude are applied, at the cost of decreasing the solution accuracy. Empirical evidence shows that such methods tend to scale nearly linearly to the system size, however, they do not exhibit provable theoretical guarantees. Similar methods have been studied based on finite-differences, which aim to approximate the Fermi distribution with polynomial expansions [145, 144]. Finally, the work of [65, 137] offer the possibility to achieve lower complexities than $O(n^{3})$ for both the eigenvalues and eigenvectors of structured matrices, by leveraging the fast multipole method. However, an end-to-end stability analysis of those algorithms remains open. Indeed, we are not aware of any end-to-end analysis with provable approximation guarantees that has lower than $O(n^{3})$ worst-case complexity for any of the aforementioned algorithms.
@@ F.3 Electron density queries
First we recall that $\epsilon\in(0,1/12)$ and note that $\|\widetilde{\boldsymbol{\mathrm{P}}}\|\leq\|\boldsymbol{\mathrm{P}}\|+\epsilon\leq 1+\epsilon.$ The quadratic form $\boldsymbol{\mathrm{x}}^{*}\widetilde{\boldsymbol{\mathrm{P}}}\boldsymbol{\mathrm{x}}$ is computed in two steps: $\boldsymbol{\mathrm{y}}\leftarrow\boldsymbol{\mathsf{fl}}(\widetilde{\boldsymbol{\mathrm{P}}}\boldsymbol{\mathrm{x}})$ and $\widetilde{n}(\boldsymbol{\mathrm{r}})=\boldsymbol{\mathsf{fl}}(\boldsymbol{\mathrm{x}}^{*}\boldsymbol{\mathrm{y}})$. From [70], Eq. (3.12) and Lemma 6.6, we know that the matrix-vector product satisfies $\boldsymbol{\mathrm{y}}=\widetilde{\boldsymbol{\mathrm{P}}}\boldsymbol{\mathrm{x}}+\boldsymbol{\mathrm{e}}$ where $\|\boldsymbol{\mathrm{e}}\|=\|\widetilde{\boldsymbol{\mathrm{P}}}\boldsymbol{\mathrm{x}}-\boldsymbol{\mathrm{y}}\|\leq\sqrt{n}\eta_{n}\|\widetilde{\boldsymbol{\mathrm{P}}}\|\|\boldsymbol{\mathrm{x}}\|$, and $\eta_{n}:=\frac{n\textbf{{u}}}{1-n\textbf{{u}}}$. For the chosen u we have $\sqrt{n}\eta_{n}\leq 2\epsilon$. From the same equations, the subsequent dot product satisfies $\widetilde{n}(\boldsymbol{\mathrm{r}})=\boldsymbol{\mathsf{fl}}(\boldsymbol{\mathrm{x}}^{*}\boldsymbol{\mathrm{y}})=\boldsymbol{\mathrm{x}}^{*}\boldsymbol{\mathrm{y}}+\epsilon$ where $|\epsilon|\leq\eta_{n}\|\boldsymbol{\mathrm{x}}\|\|\boldsymbol{\mathrm{y}}\|\leq\eta_{n}\|\boldsymbol{\mathrm{x}}\|\left(\|\widetilde{\boldsymbol{\mathrm{P}}}\boldsymbol{\mathrm{x}}\|+\|\boldsymbol{\mathrm{e}}\|\right)\leq\eta_{n}\|\widetilde{\boldsymbol{\mathrm{P}}}\|\|\boldsymbol{\mathrm{x}}\|^{2}\left(1+\sqrt{n}\eta_{n}\right)\leq 4\epsilon\|\boldsymbol{\mathrm{x}}\|^{2}$. Then $|n(\boldsymbol{\mathrm{r}})-\widetilde{n}(\boldsymbol{\mathrm{r}})|=|\boldsymbol{\mathrm{x}}^{*}\boldsymbol{\mathrm{P}}\boldsymbol{\mathrm{x}}-(\boldsymbol{\mathrm{x}}^{*}\boldsymbol{\mathrm{y}}+\epsilon)|=|\boldsymbol{\mathrm{x}}^{*}\boldsymbol{\mathrm{P}}\boldsymbol{\mathrm{x}}-\boldsymbol{\mathrm{x}}^{*}\widetilde{\boldsymbol{\mathrm{P}}}\boldsymbol{\mathrm{x}}-\boldsymbol{\mathrm{x}}^{*}\boldsymbol{\mathrm{e}}-\epsilon|\leq|\boldsymbol{\mathrm{x}}^{*}\boldsymbol{\mathrm{P}}\boldsymbol{\mathrm{x}}-\boldsymbol{\mathrm{x}}^{*}\widetilde{\boldsymbol{\mathrm{P}}}\boldsymbol{\mathrm{x}}|+|\boldsymbol{\mathrm{x}}^{*}\boldsymbol{\mathrm{e}}|+|\epsilon|\leq 2\epsilon\|\boldsymbol{\mathrm{x}}\|^{2}+2\epsilon\|\boldsymbol{\mathrm{x}}\|^{2}+4\epsilon\|\boldsymbol{\mathrm{x}}\|^{2}=8\epsilon\|\boldsymbol{\mathrm{x}}\|^{2}.$
@@
If $N$ is asymptotically larger than the system size, i.e., $N=O(\operatorname{poly}(n))$, the query complexity can be reduced by applying the Johnson-Lindenstrauss (JL) lemma [78], at the cost of a polynomial dependence on the accuracy. In brief, since the matrix $\boldsymbol{\mathrm{P}}\boldsymbol{\mathrm{S}}$ is an orthogonal projector, and therefore idempotent, it can be deduced that
@@
where $\boldsymbol{\mathrm{L}}$ is the lower triangular Cholesky factor of $\boldsymbol{\mathrm{S}}^{-1}$. Then $\|\boldsymbol{\mathrm{x}}^{*}\boldsymbol{\mathrm{P}}\boldsymbol{\mathrm{L}}^{-*}\|^{2}$ can be approximated by $\|\boldsymbol{\mathrm{x}}^{*}\boldsymbol{\mathrm{P}}\boldsymbol{\mathrm{L}}^{-*}\boldsymbol{\mathrm{T}}\|^{2}$, where $\boldsymbol{\mathrm{T}}$ is a $n\times r$ sparse random sign matrix scaled by $1/\sqrt{r}$ [1], with $r=O(\log(N)/\epsilon^{2})$ columns. The representation of $\boldsymbol{\mathrm{T}}$ requires $O(n\times r)=O(n\log(N)/\epsilon^{2})$ random bits. Alternative random matrix distributions satisfying the JL property, with varying sparsity, runtime, and dimension requirements exist in the literature [3, 82, 110]. The value of $r$ is worst-case optimal [93], however, tighter bounds have been studied in the literature [21, 126, 23, 49], thus potentially leading to more practical algorithms. The full analysis is omitted.